OpenStudy (anonymous):

Match each expression with its value. 1. −5 · −7 2. −13 − 9 3. 6 · −9 4. 18 + −4 a. −54 b. 35 c. −22 d. 14

pooja195 (pooja195):

Lets ignore the signs for right now 1) 7 x 5 = ? 2) -13-9=? 3)6 x 9 = ? 4) 18-4=?

OpenStudy (anonymous):

OMG YOUR SO SMART :D

OpenStudy (anonymous):

35=A -22=B idk

OpenStudy (anonymous):

14 is d

pooja195 (pooja195):

ok so far we have 1) A 2)B 3) we dont know yet 4) D

pooja195 (pooja195):

Whats 9 x 6 ?

OpenStudy (anonymous):

idk how c is -54 tho

OpenStudy (anonymous):

it 54 obv

OpenStudy (anonymous):

why is it negative?

OpenStudy (anonymous):

*Scratches Head*

OpenStudy (anonymous):

;O

OpenStudy (anonymous):

thanks anyways ;D mind helping me n another question?

pooja195 (pooja195):

This is why

pooja195 (pooja195):

I can try

OpenStudy (anonymous):

Which number is greater than 7.3 × 10^8?

pooja195 (pooja195):

Post the question on a new post